# Spectral Dataset of Young Type Ib Supernovae and their Time-evolution

### v1

@@
Spectral Dataset of Young Type Ib Supernovae and their Time-evolution
Due to high-cadence automated surveys, we can now detect and classify supernovae (SNe) within a few days after explosion, if not earlier. Early-time spectra of young SNe directly probe the outermost layers of the ejecta, providing insights into the extent of stripping in the progenitor star and the explosion mechanism in the case of core-collapse supernovae. However, many SNe show overlapping observational characteristics at early time, complicating the early-time classification. In this paper, we focus on the study and classification of Type Ib supernovae (SNe Ib), which are a subclass of core-collapse supernovae that lack strong hydrogen lines but show helium lines in their spectra. Here we present a spectral dataset of 8 SNe Ib, chosen to have at least 3 pre-maximum spectra, which we call early spectra. Our dataset was obtained mainly by the the Las Cumbres Observatory (LCO) and consists of a total of 82 optical photospheric spectra, including 38 early spectra. This data set increases the number of published SNe Ib with at least three early spectra by 60%. For our classification efforts, we use early spectra in addition to spectra taken around maximum light. We also convert our spectra into SN Identification (SNID) templates and make them available to the community for easier identification of young SNe Ib. Our data set increases the number of publicly available SNID templates of early spectra of SNe Ib by 43. Almost half of our sample has SN types that change over time or are different from what is listed on the Transient Name Server (TNS). We discuss the implications of our dataset and our findings for current and upcoming SN surveys and their classification efforts.
@@
Most massive stars end their life cycles as core-collapse supernovae (CCSNe), which are categorized by their spectral features: Type I supernovae (SNe) lack H lines, while Type II SNe exhibit them. Type I CCSNe are further divided into subclasses, which are collectively called Stripped Envelope Supernovae (SESNe, Clocchiatti et al. 1997; Filippenko 1997; Gal-Yam 2017; Modjaz et al. 2019), as they have been stripped of their outer layers: Type Ib SNe (SNe Ib) are characterized by the presence of strong He lines and the absence of strong H lines (though even SNe Ib may have a weak H line (Liu et al., 2016a)), and Type Ic SNe (SNe Ic) are devoid of both H and He lines. Transitional Type IIb SNe (SNe IIb) show diminishing H lines over time, but with emerging He lines akin to SNe Ib. Additionally, a subset of high-velocity SNe Ic with broad lines (SNe Ic-bl, e.g., Modjaz et al. 2016) has been interesting because they are the only kind of CCSNe observed to be connected to long-duration gamma-ray bursts (for reviews see e,g., Woosley & Bloom 2006; Modjaz 2011; Cano et al. 2017).
| SN Name | SN Typea | z | Nspecb | Date of maximum light in V-band (MJD) | Statistical Uncertainty (MJD) |
@@
b The total number of photospheric spectra for the corresponding SN included in this work.
c The earliest spectrum of SN 2022nyo, obtained on 2022-07-03 at 01:55:28 (with = -13.8 days), shows a prominent hydrogen-like feature around 6200 . This characteristic signature of a hydrogen line is further validated by the SNID results, which identify the best match with SN 2008cw (Type IIb). However, subsequent spectra acquired from 2022-07-12 at 18:29:14 (with = -4.1 days) display spectral characteristics consistent with SNe Ib.
d SN 2021ukt is a peculiar SN Ib with IIn-like features (see the section 4.3 about this SN): at early times, the TNS spectrum showed only H and no He absorption lines, thus the SN IIn classificaton above, but over times the multiple He absorption lines arise clearly, thus the SN Ib classification, while the narrow H line persisted at weaker, but varying, levels of strength.
@@
e For these 2 SNe, the LCO V-band light curves do not capture the rise of the peak. Therefore, the i-band light curve is used to calculate the date of the peak and is later converted to the date of the peak in the V-band using Table 10 of Bianco et al. (2014), assuming these SN lightcurves have the same behavior of the SESNe in Bianco et al. (2014). Conversion of the date of the peak in the i band to the date of the peak in the V band adds a systematic uncertainty of 1.5 days. We note that the additional systematic uncertainty of 1.5 days from converting from i-band to V-band is not listed in the entries above.
@@
f SNID initially classifies SN 2019odp as broad-line Type Ic (Ic-bl) based on its spectra taken before the V-band maximum light ( ¡ 0), but later classifies it as Type Ib based on the spectra taken after the V-band maximum light ( ¿ 0). See Section 4.7 for more information.
@@
Despite the understanding that SESNe result from the core collapse of massive stars, the diverse range of observed explosion properties poses a challenge in classifying SESNe into distinct types. As a result, it remains unclear if the classification types are distinct or share overlapping characteristics, and what properties of the progenitors or of the explosion cause this diversity in the observed SNe properties. Addressing these unresolved questions involves two approaches: (i) analyzing data of large samples to identify trends in different supernova (SN) types and (ii) studying data of SNe well before maximum light to directly probe the outermost layers of the progenitor envelope, as the photosphere lies in the outermost layers of the ejecta during this early phase and recedes into deeper layers over time.
@@
Traditionally, individual SNe have been thoroughly studied, such as Type IIb SN 1993J (Filippenko et al. 1993; Matheson et al. 2000b; Matheson et al. 2000a) and Type Ic SN 1994I (Filippenko et al. 1995; Richmond et al. 1996). The prevalence of population studies has increased in recent years. Statistical comparisons of SESNe spectra by Matheson et al. (2000a) revealed heterogeneous properties, while subsequent studies by Fremling et al. (2018), Prentice et al. (2019), Liu et al. (2016a), and Holmbo et al. (2023) explored the spectral properties of different types of SESNe. To quantify and visualize the continuum of SESNe classification, Williamson et al. (2019a) presented a new quantitative and data-driven classification technique utilizing the machine learning tool of Support Vector Machine (SVM) after applying Principal Component Analysis (PCA) to the photospheric spectra of SESNe. They found that the SESNe types, in particular SNe Ib, are most distinguishable in later phase ranges, particularly 10 5 days and 15 5 days relative to V-band maximum.
@@
However, these studies lacked early spectra in their sample because they included data from earlier surveys, which had low cadences. Advancements in high-cadence wide-field surveys now enable prompt detection and observation of SNe. Through these early time observations, though, we are discovering an increasing number of SNe with evolving properties, such as SNe that change spectroscopic type throughout their evolution. These SNe pose a challenge to our classification efforts. The transitional SN 2022crv (Dong et al., 2023a; Gangopadhyay et al., 2023) is an example of an evolving SESN.
@@
Now is the perfect time to start conducting comprehensive population studies of young SESNe, in particular SNe Ib. This work constitutes the first step in these population studies, as it publishes optical spectra and spectral templates of young SNe Ib only days after explosion, well before maximum light. Studying the spectra of young SNe Ib over time will provide valuable insights into when the tell-tale He lines will appear. Consequently, it will offer significant constraints on 56Ni mixing, as the He lines result from non-thermal excitation induced by fast electrons that are accelerated by energetic MeV photons emitted during Ni decay (Lucy, 1991). Tracking the appearance of the He lines helps with inferring the distribution of Ni-mixing further out in the ejecta. In turn, the exact details of the Ni-mixing can constrain explosion models, which, for objects where the He lines are visible shortly after explosion, would require the release of Ni close to the outermost He layer, potentially indicating large scale turbulence (e.g., Hammer et al. 2010). Furthermore, analyzing early spectra will inform the most appropriate timing to schedule spectroscopy to classify an SN Ib, which is essential for the new era of big-data transient science, in which thousands of SNe will be photometrically discovered every night with the Vera C. Rubin Observatory as part of the Legacy Survey of Space and Time (LSST; Ivezić et al. 2019).
@@
In Section 2, we present a dataset of 82 photospheric spectra of 8 different SNe Ib, including 38 spectra taken before maximum light. All spectra will be made public via the Weizmann Interactive Supernova Data REPository (WISeREP222https://www.wiserep.org/; Yaron & Gal-Yam 2012). We discuss our spectrum phase calculation and the method for calculating date of maximum light in Section 3. In Section 4, we discuss the specifics of the Supernova Identification (SNID) code (Blondin & Tonry, 2007) classification of each SN in the sample. In Section 5, we present our procedure for creating spectral templates for SNID for the SNe in our sample using our data as well as prior published data. Our SNID templates can be downloaded via our METAL GitHub repository333https://github.com/metal-sn/SESNtemple/tree/master, formerly known as the SNYU github page.
@@ 2 Spectroscopic Observations and Data Reduction
Our sample consists of 8 SNe Ib with data taken between 2016 and 2023 mainly by the Las Cumbres Observatory (LCO) (Brown et al., 2013) as part of the Global Supernova Project (GSP) (Howell, 2024). These 8 SNe Ib were selected from the GSP sample because a) they had observed photometry that covers the lightcurve peak in at least one band, so spectral phases could be calculated relative to the date of maximum (see Section 3 for details), and b) they had unpublished spectra, with at least three spectra before the date of V-band maximum light (t), such that we could study their early-time behavior. Additional data were taken as part of the LCO Supernova Key Project (Howell, 2017) for SN 2016bau and by GSP members with other observatories.
@@
Thus, we present a total of 82 photospheric (t ¡ 60 days) optical spectra of 8 SNe Ib. We exclude nebular-phase spectra for this study since we focus on photospheric-phase classification. Our data set includes a total of 38 spectra taken before the V-band maximum light, which we refer to as early spectra, increasing the number of published SNe Ib with at least 3 early spectra by 60 and the number of publicly available early spectra of SNe Ib that satisfy the same criterion of at least three early spectra by 37. Table 1 shows a summary of our data set. Figure 1 shows the spectral time series of the 8 SNe Ib in our sample.
@@
Almost all (80 out of 82) optical spectra were taken using the FLOYDS spectrographs on the 2m telescopes at Siding Spring Observatory (COJ 2m), Australia, and Haleakala (OGG 2m), Hawaii. A 2 arcsec slit was placed on the SN along the parallactic angle (Filippenko, 1982). The spectra are extracted from the raw images with wavelength scale and flux calibration applied by the FLOYDS pipeline 444https://lco.global/documentation/data/floyds-pipeline/, following standard procedures (Valenti et al., 2014). One optical spectrum was taken with the Southern African Large Telescope (SALT) using the Robert Stobe Spectrograph (RSS; Smi 2006) through Rutgers University program 2022-1-MLT-004 (PI: Jha). The observations were taken with the PG0900 grating and 1.5″wide longslit, giving a spectral resolution . The data were reduced using a custom pipeline based on standard Pyraf spectral reduction routines and the PySALT package (Crawford et al., 2010). Additionally, one spectrum was obtained with the Goodman High Throughput Spectrograph on the Southern Astrophysical Research (SOAR) 4.1m telescope. The spectrum was taken using the blue camera, covering the wavelength range 3500 – 7000 Å, with a 1″slit and 400 mm-1 line grating. We used the Goodman Spectroscopic Data Reduction Pipeline to perform bias and flat corrections, cosmic-ray removal, and wavelength calibration. Fluxes were calibrated to a standard star observed on the same night with the same instrumental setup.
The details of the observation log are presented in Appendix A Table LABEL:tab:_obs_log. All the spectra presented in this work will be made available through WISeREP.
@@ 3 Spectrum Phases and V-band Maximum Date Calculation
Determining the phase of each spectrum, or the epoch relative to maximum light, is essential for conducting a systematic analysis of different SN properties. Our phase calculations are based on the V-band maximum date (), ensuring consistent tracking of the appearance of He lines across our sample. This approach also facilitates consistent comparisons with other population studies, such as those conducted by Liu et al. (2016a) and Holmbo et al. (2023), which similarly adopt the time of the V-band maximum date as their reference point for phase calculation.
@@
While all SNe in our sample have V-band light curves taken by LCO, 2 SNe lack a clear peak in the V-band light curve. For these two cases, namely SNe 2021hen and 2021ukt, we utilized the LCO i-band light curve to determine the maximum date in the i-band, as redder bands exhibit maxima at later times. We then convert the determined maximum date in the i-band to the maximum date in the V-band using Table 10 from Bianco et al. (2014), assuming these SN lightcurves have the same behavior as the SESNe in Bianco et al. (2014), which for SN 2021ukt is a large caveat (see section 4.3). This step adds a systematic uncertainty of 1.5 days. To measure the date of maximum light for the SNe in our sample, we follow Bianco et al. (2014) and use a Monte Carlo (MC) simulation process to fit a quadratic around the peak of the light curve for a given filter. The adopted date of maximum light is the mean of the maximum epochs across the MC realizations, and the error associated with the date of maximum light is the corresponding standard deviation. Table 1 shows the dates of maximum in the V band and their associated uncertainties that we calculated for each of the SNe in our data set.
We compare our calculated maximum dates for two SNe in our sample with previously published results and find general agreement. For SN 2019odp, we find a V-band maximum date of Modified Julian Date (MJD) 58735.19 0.16, which is consistent with the findings of Schweyer et al. (2023), who reported a maximum date of MJD 58734 1 day in the g-band. For SN 2016bau, Aryan et al. (2021) found the maximum date in the B-band to be MJD 57477.37 1.99 days. For comparison, we converted our calculated V-band maximum date to the B-band equivalent by using Table 10 from Bianco et al. (2014), resulting in MJD 57475.87 0.69 days, with an additional systematic uncertainty of 1.3 days. Thus, our measured date of max for SN 2016bau is consistent within the uncertainties with the published determination by Aryan et al. (2021).

### v2

# Spectral dataset of young type Ib supernovae and their time evolution

@@
11email: pqt7tv@virginia.edu, vru7qe@virginia.edu 22institutetext: Las Cumbres Observatory, 6740 Cortona Drive, Suite 102, Goleta, CA 93117-5575, USA 33institutetext: Department of Physics, University of California, Santa Barbara, CA 93106-9530, USA 44institutetext: School of Physics and Astronomy, Tel Aviv University, Tel Aviv 69978, Israel 55institutetext: Center for Astrophysics, Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138-1516, USA 66institutetext: The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, USA 77institutetext: Steward Observatory, University of Arizona, 933 North Cherry Avenue, Tucson, AZ 85721-0065, USA 88institutetext: Department of Astronomy & Astrophysics, University of California, San Diego, 9500 Gilman Drive, MC 0424, La Jolla, CA 92093-0424, USA 99institutetext: Department of Physics and Astronomy, Rutgers, The State University of New Jersey, 136 Frelinghuysen Road, Piscataway, NJ 08854-8019, USA
Spectral dataset of young type Ib supernovae and their time evolution
Due to high-cadence automated surveys, we can now detect and classify supernovae (SNe) within a few days after explosion, if not earlier. Early-time spectra of young SNe directly probe the outermost layers of the ejecta, providing insights into the extent of stripping in the progenitor star and the explosion mechanism in the case of core-collapse supernovae. However, many SNe show overlapping observational characteristics at early times, complicating the early-time classification. In this paper, we focus on the study and classification of type Ib supernovae (SNe Ib), which are a subclass of core-collapse SNe that lack strong hydrogen lines but show helium lines in their spectra. Here we present a spectral dataset of eight SNe Ib, chosen to have at least three pre-maximum spectra, which we call early spectra. Our dataset was obtained mainly by the Las Cumbres Observatory (LCO) and it consists of a total of 82 optical photospheric spectra, including 38 early spectra. This dataset increases the number of published SNe Ib with at least three early spectra by 60%. For our classification efforts, we used early spectra in addition to spectra taken around maximum light. We also converted our spectra into SN IDentification (SNID) templates and make them available to the community for easier identification of young SNe Ib. Our dataset increases the number of publicly available SNID templates of early spectra of SNe Ib by 43. Half of our sample has SN types that change over time or are different from what is listed on the Transient Name Server (TNS). We discuss the implications of our dataset and our findings for current and upcoming SN surveys and their classification efforts.
@@
Most massive stars end their life cycles as core-collapse supernovae (CCSNe), which are categorized by their spectral features: type I supernovae (SNe) lack H lines, while type II SNe exhibit them. type I CCSNe are further divided into subclasses, which are collectively called stripped envelope supernovae (SESNe, Clocchiatti et al. 1997; Filippenko 1997; Gal-Yam 2017; Modjaz et al. 2019), as they have been stripped of their outer layers: type Ib SNe (SNe Ib) are characterized by the presence of strong He lines and the absence of strong H lines (though even SNe Ib may have a weak H line Liu et al. 2016), and type Ic SNe (SNe Ic) are devoid of both H and He lines. Transitional type IIb SNe (SNe IIb) show diminishing H lines over time, but with emerging He lines akin to SNe Ib. Additionally, a subset of high-velocity SNe Ic with broad lines (SNe Ic-bl, e.g., Modjaz et al. 2016) has been interesting because they are the only kind of CCSNe observed to be connected to long-duration gamma-ray bursts (for reviews, see, e.g., Woosley & Bloom 2006; Modjaz 2011; Cano et al. 2017).
| SN name | SN typea | zb | Nspecc | Date of maximum light in V-band (MJD) | Statistical uncertainty (MJD) |
@@
b Redshift values are obtained from TNS, except for SN 2022nyo, for which the redshift was assumed to be the redshift of its host galaxy available on NASA/IPAC Extragalactic Database (NED; Helou et al. 1991).
@@
c The total number of photospheric spectra for the corresponding SN included in this work.
@@
d For these 2 SNe, the LCO V-band light curves do not capture the rise to the peak. Therefore, the i-band light curve is used to calculate the date of the peak and is later converted to the date of the peak in the V-band using Table 10 of Bianco et al. (2014), assuming these SN lightcurves have the same behavior of the SESNe in Bianco et al. (2014). Conversion of the date of the peak in the i band to the date of the peak in the V band adds a systematic uncertainty of 1.5 days. We note that the additional systematic uncertainty of 1.5 days from converting from i-band to V-band is not listed in the entries above. Additionally, using the conversion relationship between the date of maximum in the V-band and i-band from Rodríguez et al. (2023) (see Section 3), we obtain = 59442.06 1.63 and 59312.13 1.62 for SNe 2021ukt and 2021hen respectively.
e The earliest spectrum of SN 2022nyo, obtained on UT 2022-07-03 01:55:28 (with = 13.8 days), shows a prominent hydrogen-like feature around 6200 . This characteristic signature of a hydrogen line is further validated by the SNID results, which identify the best match with SN 2008cw (type IIb). However, subsequent spectra acquired on UT 2022-07-12 18:29:14 (with = 4.1 days) display spectral characteristics consistent with SNe Ib (see Section 4.2 for further details).
f SN 2021ukt is a peculiar SN Ib with IIn-like features (see the section 4.3 about this SN): at early times, the TNS spectrum showed only H and no He absorption lines, thus the SN IIn classification above, but over times the multiple He absorption lines arise clearly, thus the SN Ib classification, while the narrow H line persisted at weaker, but varying, levels of strength.
@@
g SNID initially classifies SN 2019odp as broad-line type Ic (Ic-bl) based on its spectra taken before the V-band maximum light ( ¡ 0), but later classifies it as type Ib based on the spectra taken after the V-band maximum light ( ¿ 0). See Section 4.7 for more information.
@@
Despite the understanding that SESNe result from the core collapse of massive stars, the diverse range of observed explosion properties poses a challenge in classifying SESNe into distinct types. As a result, it remains unclear if the classification types are distinct or share overlapping characteristics, and what properties of the progenitors or of the explosion cause this diversity in the observed SNe properties. Addressing these unresolved questions involves two approaches: (i) analyzing data of large samples to identify trends in different SN types and (ii) studying data of SNe well before maximum light to directly probe the outermost layers of the progenitor envelope, as the photosphere lies in the outermost layers of the ejecta during this early phase and recedes into deeper layers over time.
@@
Traditionally, individual SNe have been thoroughly studied, such as type IIb SN 1993J (Filippenko et al. 1993; Matheson et al. 2000a, b) and type Ic SN 1994I (Filippenko et al. 1995; Richmond et al. 1996). The prevalence of population studies has increased in recent years. Statistical comparisons of SESNe spectra by Matheson et al. (2001) revealed heterogeneous properties, while subsequent studies by Fremling et al. (2018), Prentice et al. (2019), Liu et al. (2016), and Holmbo et al. (2023) explored the spectral properties of different types of SESNe. To quantify and visualize the continuum of SESNe classification, Williamson et al. (2019) presented a new quantitative and data-driven classification technique utilizing the machine learning tool of the support vector machine (SVM) after applying principal component analysis (PCA) to the photospheric spectra of SESNe. They find that the SESNe types, in particular SNe Ib, are most distinguishable in later phase ranges, particularly 10 5 days and 15 5 days relative to the V-band maximum.
@@
However, these studies lack early spectra in their sample because they include data from earlier surveys that have low cadences. Advancements in high-cadence wide-field surveys now enable prompt detections and observations of SNe to be made. Through these early time observations, though, we are discovering an increasing number of SNe with evolving properties, such as SNe that change their spectroscopic type throughout their evolution. These SNe pose a challenge to our classification efforts. The transitional SN 2022crv (Dong et al., 2024; Gangopadhyay et al., 2023) is an example of an evolving SESN, transitioning from type IIb to type Ib in the early time.
@@
Now is the perfect time to start conducting comprehensive population studies of young SESNe, in particular SNe Ib. This work constitutes the first step in these population studies, as it publishes optical spectra and spectral templates of young SNe Ib only days after explosion, well before maximum light. Studying the spectra of young SNe Ib over time will provide valuable insights into when the tell-tale He lines will appear. Consequently, it will offer significant constraints on 56Ni mixing, as the He lines result from nonthermal excitation induced by fast electrons that are accelerated by energetic megaelectron-volt photons emitted during Ni decay (Lucy, 1991). Tracking the appearance of the He lines helps with inferring the distribution of Ni-mixing further out in the ejecta. In turn, the exact details of the Ni mixing can constrain explosion models, which, for objects where the He lines are visible shortly after explosion, would require the release of Ni close to the outermost He layer, potentially indicating large-scale turbulence (e.g., Hammer et al. 2010). Furthermore, analyzing early spectra will inform the most appropriate timing to schedule spectroscopy to classify an SN Ib, which is essential for the new era of big-data transient science, in which thousands of SNe will be photometrically discovered every night with the Vera C. Rubin Observatory as part of the Legacy Survey of Space and Time (LSST; Ivezić et al. 2019).
@@
In Section 2, we present a dataset of 82 photospheric spectra of eight different SNe Ib, including 38 spectra taken before maximum light. All spectra will be made public via the Weizmann Interactive Supernova Data REPository (WISeREP222https://www.wiserep.org/; Yaron & Gal-Yam 2012). We discuss our spectrum phase calculation and the method for calculating date of maximum light in Section 3. In Section 4, we discuss the specifics of the SuperNova IDentification (SNID) code (Blondin & Tonry, 2007) classification of each SN in the sample. In Section 5, we present our procedure for creating spectral templates for SNID for the SNe in our sample using our data as well as prior published data. Our SNID templates can be downloaded via our METAL GitHub repository333https://github.com/metal-sn/SESNtemple/tree/master, formerly known as the SNYU github page.
@@ 2 Spectroscopic observations and data reduction
Our sample consists of eight SNe Ib with data taken between 2016 and 2023 mainly by the Las Cumbres Observatory (LCO) (Brown et al., 2013) as part of the Global Supernova Project (GSP) (Howell, 2019, 2024). These eight SNe Ib were selected from the GSP sample because a) they had observed photometry that covers the lightcurve peak in at least one band, so spectral phases could be calculated relative to the date of maximum (see Section 3 for details), and b) they had unpublished spectra, with at least three spectra before the date of V-band maximum light (t), such that we could study their early-time behavior. Additional data were taken as part of the LCO Supernova Key Project (Howell, 2017) for SN 2016bau and by GSP members with other observatories.
@@
Thus, we present a total of 82 photospheric (t ¡ 60 days) optical spectra of eight SNe Ib. We exclude nebular-phase spectra for this study since we focus on photospheric-phase classification. Our dataset includes a total of 38 spectra taken before the V-band maximum light, which we refer to as early spectra, increasing the number of published SNe Ib with at least three early spectra by 60 and the number of publicly available early spectra of SNe Ib that satisfy the same criterion of at least three early spectra by 37. Table 1 shows a summary of our dataset. Figure 1 shows an illustrative example of the spectral time series for two SNe Ib from our sample, while Figure 3 presents the complete spectral time series for all eight SNe Ib in our study.
@@
Almost all (80 out of 82) optical spectra were taken using the FLOYDS spectrographs on the 2m telescopes at Siding Spring Observatory (COJ 2m), Australia, and Haleakala (OGG 2m), Hawaii. A 2″slit was placed on the SN along the parallactic angle (Filippenko, 1982). The spectra are extracted from the raw images with wavelength scale and flux calibration applied by the FLOYDS pipeline 444https://lco.global/documentation/data/floyds-pipeline/, following standard procedures (Valenti et al., 2014). One optical spectrum was taken with the Southern African Large Telescope (SALT) using the Robert Stobie Spectrograph (RSS; Smi 2006) through Rutgers University program 2022-1-MLT-004 (PI: Jha). The observations were taken with the PG0900 grating and 1.5″wide longslit, giving a spectral resolution . The data were reduced using a custom pipeline based on standard Pyraf spectral reduction routines and the PySALT package (Crawford et al., 2010). Additionally, one spectrum was obtained with the Goodman High Throughput Spectrograph on the Southern Astrophysical Research (SOAR) 4.1m telescope. The spectrum was taken using the blue camera, covering the wavelength range 3500 – 7000 Å, with a 1″slit and 400 mm-1 line grating. We used the Goodman Spectroscopic Data Reduction Pipeline to perform bias and flat corrections, cosmic-ray removal, and wavelength calibration. Fluxes were calibrated to a standard star observed on the same night with the same instrumental setup. The details of the observation log are presented in Appendix B Table LABEL:tab:_obs_log. All the spectra presented in this work will be made available through WISeREP.
@@
| SN name | Our calculated (MJD) | Our calculated (MJD) | from published work (MJD) | from published work (MJD) | Reference | Method of conversion a |
| SN 2020hvp | 58975.95 | 0.46 | 58975.7 | 0.4 | Rodríguez et al. (2023) | - |
| SN 2019odp | 58735.19 | 0.16 | 58735.2b | 1.1b | Schweyer et al. (2024) | R1 |
| 58736.1 | 0.6 | Rodríguez et al. (2023) | - | |||
| SN 2016bau | 57478.17 | 0.69 | 57479.37c | 2.23c | Aryan et al. (2021) | R1 |
| 57479.67c | 1.99c, d | Aryan et al. (2021) | B10 | |||
| 57477.7 | 0.2 | Rodríguez et al. (2023) | - |
b Converted from .
c Converted from .
d An additional systematic uncertainty of 1.3 days, resulting from the conversion between B-band and V-band, is not included in the reported uncertainty above.
@@ 3 Spectrum phases and V-band maximum date calculation
Determining the phase of each spectrum, or the epoch relative to maximum light, is essential for conducting a systematic analysis of different SN properties. Our phase calculations are based on the V-band maximum date (), ensuring consistent tracking of the appearance of He lines across our sample. This approach also facilitates consistent comparisons with other population studies, such as those conducted by Liu et al. (2016) and Holmbo et al. (2023), which similarly adopt the time of the V-band maximum date as their reference point for phase calculation.
@@
While all SNe in our sample have V-band light curves taken by LCO, 2 SNe lack a clear peak in the V-band light curve. For these two cases, namely SNe 2021hen and 2021ukt, we utilized the LCO i-band light curve to determine the maximum date in the i-band (), as redder bands exhibit maxima at later times. We then convert the determined maximum date in the i-band to the maximum date in the V-band using Table 10 from Bianco et al. (2014), assuming these SN lightcurves have the same behavior as the SESNe in Bianco et al. (2014), which for SN 2021ukt is a large caveat (see section 4.3). This step adds a systematic uncertainty of 1.5 days. Alternatively, we could use the average values between and the epochs of maximum light in other optical bands for SNe Ib, as outlined in Table 1 of Rodríguez et al. (2023). This approach yields a conversion equation for the date of maximum light between the V-band and i-band, given by = ( + 3.1) 1.6. Using this conversion relationship, the maximum dates for SNe 2021hen and 2021ukt are provided in the footnote of Table 1. However, we opt to use the converted dates from Bianco et al. (2014) for the phase calculations of our spectra to be consistent. See below for more details.
We perform the maximum light date calculations using the light curves after their preliminary reductions; as such, we do not publish the full photometry in this paper. To measure the date of maximum light for the SNe in our sample, we follow Bianco et al. (2014) and use a Monte Carlo (MC) simulation process to fit a quadratic around the peak of the light curve for a given filter. The adopted date of maximum light is the mean of the maximum epochs across the MC realizations, and the error associated with the date of maximum light is the corresponding standard deviation. Table 1 shows the dates of maximum in the V band and their associated uncertainties that we calculated for each of the SNe in our dataset.
We compare our calculated maximum dates for several SNe in our sample with previously published results and find agreement in all cases. Table 2 presents a comparison between our calculated maximum dates and those from previously published results. For SN 2019odp, we find a V-band maximum date of Modified Julian Date (MJD) 58735.19 0.16. In comparison, Schweyer et al. (2023) reported a maximum date of MJD 58734 1 day in the g-band (), calculated with their own data. To convert the date of maximum in the g-band to the V-band, we use the average value between and from Table 1 of Rodríguez et al. (2023), which gives the relationship specifically for SNe Ib: = ( + 1.2) 0.5. This conversion yields = 58735.2 1.1 for SN 2019odp, consistent with our result. Furthermore, Rodríguez et al. (2023) determined the V-band maximum for SN 2019odp as MJD 58736.1 0.6, based on data from both the literature and the Zwicky Transient Facility (ZTF; Graham et al. 2019), further supporting the consistency of our measurements. For SN 2020hvp, our determined V-band maximum of MJD 58975.95 0.46 is in agreement with Rodríguez et al. (2023), who report MJD 58975.7 0.4, indicating consistency within the uncertainties. For SN 2016bau, Aryan et al. (2021) found that the maximum date in the B band () was MJD 57477.37 1.99 days, calculated from their own data. For comparison, we converted the to our calculated of MJD 57478.17 0.69 by using Table 10 from Bianco et al. (2014), resulting in converted MJD 57479.67 1.99 days, with an additional systematic uncertainty of 1.3 days. One caveat here is that this conversion relationship from Bianco et al. (2014) is calculated for different SESNe subtypes, and we assume that the light curves of the SNe Ib in our sample behave similarly to those in their sample. But to further verify the conversion, we refer to the relationship reported in Table 1 of Rodríguez et al. (2023), as this conversion is specifically calculated for SNe Ib. According to Rodríguez et al. (2023), the conversion relationship is given by = ( + 2.0) 1.0, which yields = 57479.37 2.23, consistent with our result. Our result for SN 2016bau is also in agreement with Rodríguez et al. (2023), who report a V-band maximum of MJD 57477.7 0.2.